# The Unpaid Toll: Quantifying the Public Health Impact of AI

### v1

@@ Abstract
The surging demand for AI has led to a rapid expansion of energy-intensive data centers, impacting the environment through escalating carbon emissions and water consumption. While significant attention has been paid to AI’s growing environmental footprint, the public health burden, a hidden toll of AI, has been largely overlooked. Specifically, AI’s lifecycle, from chip manufacturing to data center operation, significantly degrades air quality through emissions of criteria air pollutants such as fine particulate matter, substantially impacting public health. This paper introduces a methodology to model pollutant emissions across AI’s lifecycle, quantifying the public health impacts. Our findings reveal that training an AI model of the Llama-3.1 scale can produce air pollutants equivalent to more than 10,000 round trips by car between Los Angeles and New York City. The total public health burden of U.S. data centers in 2030 is valued at up to more than \$20 billion per year, double that of U.S. coal-based steelmaking and comparable to that of on-road emissions of California. Further, the public health costs unevenly impact economically-disadvantaged communities, where the per-household health burden could be 200x more than that in less-impacted communities. We recommend adopting a standard reporting protocol for criteria air pollutants and the public health costs of AI, paying attention to all impacted communities, and implementing health-informed AI to mitigate adverse effects while promoting public health equity.
@@ 1 Introduction
The rise of artificial intelligence (AI) has numerous potentials to play a transformative role in addressing grand societal challenges, including air quality and public health [1, 2]. For example, by integrating multimodal data from various sources, AI can provide effective tools and actionable insights for pandemic preparedness, disease prevention, healthcare optimization, and air quality management [3, 1]. However, the surging demand for AI — particularly generative AI, as exemplified by the recent popularity of large language models (LLMs) — has driven a rapid increase in computational needs, fueling the unprecedented expansion of energy-intensive AI data centers. According to McKinsey projections, under a medium-growth scenario [4], the U.S. data centers are anticipated to account for 11.7% of national electricity consumption in 2030, a substantial increase from their current share of less than 4% in 2023.
@@
The growing electricity demand of AI data centers has not only created significant stress on power grid stability [5, 6], but also increasingly impacts the environment through escalating carbon emissions [7, 8] and water consumption [9]. These environmental impacts are driven primarily by the “expansion of AI products and services,” as recently acknowledged by Google in its latest sustainability report [10]. To mitigate the challenges posed to both power grids and the environment, a range of strategies have been explored, including grid-integrated data centers [6, 11], energy-efficient hardware and software [12, 13, 14], and the adoption of carbon-aware and water-efficient computing practices [15, 9, 16, 17], among others.
The hidden toll of AI.
@@
While the environmental footprint of AI has garnered attention, the public health burden, a hidden toll of AI, has been largely overlooked. Across its entire lifecycle — from chip manufacturing to data center operation — AI contributes substantially to air quality degradation and public health costs through the emission of various criteria air pollutants. These include fine particulate matter (PM2.5, particles measuring 2.5 micrometers or smaller in diameter that can penetrate deep into lungs and cause serious health effects), sulfur dioxide (SO2), and nitrogen dioxide (NO2). Concretely, the AI hardware manufacturing process [18], electricity generation from fossil fuels to power AI data centers, and the maintenance and usage of diesel backup generators to ensure continuous AI data center operation all produce significant amounts of criteria air pollutants. Moreover, the distinct spatial-temporal heterogeneities of emission sources suggest that focusing solely on reducing AI’s carbon footprints may not minimize its emissions of criteria air pollutants or the resulting public health impacts (Section 5).
@@
Exposure to criteria air pollutants is directly and causally linked to various adverse health outcomes, [^2] While we focus on public health, we note that the impacts of criteria air pollutants extend beyond humans and include harms to environmentally sensitive areas, such as some national parks and wilderness areas which, classified as “Class 1 areas” under the Clean Air Act, require special air protection [19]. including premature mortality, lung cancer, asthma, heart attacks, cardiovascular diseases, strokes, and even cognitive decline, especially for the elderly and vulnerable individuals with pre-existing conditions [20, 21, 22, 23]. Moreover, even short-term (hours to days) PM2.5 exposure is harmful and deadly, accounting for approximately 1 million premature deaths per year from 2000 to 2019 and representing 2% of total global deaths [24].
@@
While the U.S. has generally better air quality than many other countries, 4 in 10 people in the U.S. still live with unhealthy levels of air pollution, according to the “State of the Air 2024” report published by the American Lung Association [28]. In 2019 (the latest year of data provided by the World Health Organization, or WHO, as of November 2024), an estimate of 93,886 deaths in the U.S. were attributed to ambient air pollution [29]. In fact, even compliance with the U.S. Environmental Protection Agency (EPA) air quality standards does not necessarily guarantee healthy air that meets the WHO guidelines. Concretely, the EPA’s recently tightened primary standard for PM2.5 sets an annual average limit of 9 $\mu g/m^{3}$, considerably higher than the WHO’s recommended level of 5 $\mu g/m^{3}$ [30, 31]. In addition, the EPA projects that 53 U.S. counties, including 23 in the most populous state of California, would fail to meet the revised national annual PM2.5 standard in 2032 [32].
Further, criteria air pollutants are not confined to the immediate vicinity of their emission sources; they can travel hundreds of miles through a dispersion process (i.e., cross-state air pollution) [33, 34], impacting public health across vast regions — pollutants from the 2024 Canadian wildfires significantly degraded air quality across much of the U.S. and reached as far as Mexico and Europe [35].
@@
Importantly, along with transportation and industrial activities, electricity generation is a major contributor to ambient air pollution with substantial public health impacts [26, 36, 37]. For example, a recent study [38] shows that, between 1999 and 2020, a total of 460,000 excess deaths were attributed to PM2.5 generated by coal-fired power plants alone in the U.S. As highlighted by the U.S. EPA [36], despite years of progress, “fossil fuel-based power plants remain a leading source of air, water, and land pollution that affects communities nationwide.” Moreover, according to the U.S. Energy Information Administration (EIA) projection [39], the coal consumption by the electricity sector in 2050 will still be about 30% of the 2024 level in the baseline reference case, and the number will exceed 50% in the high zero-carbon technology cost case. Indeed, the growing energy demands of AI are already delaying the decommissioning of coal-fired power plants and increasing fossil-fuel plants in the U.S. as well as around the world [6, 40, 41].
The public health outcomes of AI due to its emission of criteria air pollutants lead to various losses, such as hospitalizations, medication usage, emergency room visits, school loss days, and lost workdays. Moreover, these losses can be further quantified in economic costs based on epidemiology and economics research for the corresponding health endpoints [42, 22]. In contrast, the environmental impacts of AI, e.g., carbon emission from fossil fuels and water consumption for data center cooling, often do not cause the same immediate health impacts. For instance, while anthropogenic carbon emissions could also pose risks to public health, such impacts are often second- or third-order effects through long-term climate change which can then threaten the human well-being by affecting the food people eat and facilitating the spreading of pests, among others [43]. Nonetheless, despite their immediate and tangible impacts on public health, the criteria air pollutants of AI have remained under the radar, entirely omitted from today’s AI risk assessments and sustainability reports [44, 45, 10].
Quantifying the public health costs of AI.
@@
In this paper, we uncover and quantify the hidden public health impacts of AI. We introduce a general methodology to model the emission of criteria air pollutants associated with AI tasks across three distinct scopes: emissions from the maintenance and operation of backup generators (Scope 1), emissions from fossil fuel combustion for electricity generation (Scope 2), and emissions resulting from the manufacturing of server hardware (Scope 3). Then, we analyze the dispersion of criteria air pollutants and the resulting public health impacts across different regions.
@@
Our main results (Section 4) focus on the scope-2 health impacts of U.S. data centers and, specifically, LLM training. [^3] Our study focuses on the 48 contiguous U.S. states plus Washington D.C. because the EPA data does not include other regions [46]. If located in countries with higher population densities or less strict air quality standards, the same AI task and data centers would likely contribute to significantly more deaths and other adverse health effects. We recommend further research on the public health impact of AI outside the U.S. Using the reduced-complexity modeling tool COBRA (CO-Benefits Risk Assessment) provided by the EPA [46], our screening analysis demonstrates that driven by the growing demand for AI, the U.S. data centers could contribute to, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths in 2030, exceeding 1/3 of asthma deaths in the U.S. each year [47]. The overall public health costs could reach more than \$20 billion, double that of the U.S. coal-based steelmaking industry [48], and rival or even top those of on-road emissions of the largest U.S. states such as California with $\sim$35 million registered vehicles [49]. Moreover, depending on the location, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a car for more than 10,000 round trips between Los Angeles and New York City (LA-NYC), resulting in a health cost that even exceeds 120% of the training electricity cost.
Critically, the health costs are unevenly distributed across counties and communities, disproportionately affecting low-income counties (e.g., Meigs County, Ohio) where the per-household health burden could be equivalent to nearly 8 months of electricity bills and more than 200x compared to that in other counties.
@@
In addition, to highlight the importance of scope-1 and scope-3 health impacts, we consider data center backup generators in Virginia (Scope 1) and semiconductor manufacturing plants in Arizona and Ohio (Scope 3). Our analysis shows that, assuming the actual emissions are only 10% of the permitted level, the data center backup generators registered in Virginia (mostly in Loudoun, Prince William, and Fairfax) could already cause 14,000 asthma symptom cases among other health outcomes and a total public health burden of \$220-300 million per year, impacting residents in multiple surrounding states and as far as Florida (Section 2.2.1). If these data centers emit air pollutants at the maximum permitted level, the total public health cost will become 10-fold and reach \$2.2-3.0 billion per year. The scope-3 health impact of AI is also substantial. For example, just a single semiconductor facility in Arizona can cause an annual public health cost of \$26-39 million, with \$14-21 million attributed to the facility’s on-site emissions of criteria air pollutants (Section 2.2.2). Furthermore, relocating the same facility to a planned site in Ohio could almost quadruple the public health cost to \$94-156 million, with \$23-36 million resulting from on-site emissions.
@@
Finally, we provide recommendations to address the increasing public health impact of AI (Section 5). Specifically, we recommend technology companies adopt a standard reporting protocol for criteria air pollutants and public health impacts in their AI model cards and sustainability reports, implement health-informed AI to proactively minimize the adverse health effects of AI data centers, pay attention to all impacted communities, and prioritize reducing the health impact on disadvantaged communities to promote public health equity.
@@
To summarize, our study sheds light on and quantifies the overlooked public health impact of AI. It can inform the public, policymakers, and technology companies in conducting a more comprehensive cost-benefit analysis. We also urge further research to comprehensively address the public health implications when developing powerful and truly responsible AI in the future, ensuring that the growth of AI does not exacerbate the health burden or outweigh the potential benefits AI can provide to improve public health.
2 Background on the Air Quality Impact of AI
This section presents an overview of AI’s impact on air quality and contribution to criteria air pollutants throughout its lifecycle, beginning with background on criteria air pollutants and U.S. air quality policies.
2.1 Criteria Air Pollutants
@@
While CO2 is broadly classified by the EPA as an air pollutant following the U.S. Supreme Court ruling in 2007 [53] and contributes to long-term climate change, it often does not cause the same immediate health impacts as criteria pollutants. In the U.S., CO2 and other greenhouse gases are subject to different EPA regulations from those for criteria air pollutants. Thus, for the sake of presentation in this paper, we use “air pollutants” to solely refer to criteria air pollutants wherever applicable.
2.2 AI’s Contribution to Air Pollutants
To understand the impact of AI on air quality, we focus on the three scopes over which AI contributes to criteria air pollutants as well as other toxic materials. The scoping definition in this paper parallels the well-established greenhouse gas protocol [54].
2.2.1 Scope 1
@@
The scope-1 public health impact of AI primarily comes from the emission of operating on-site backup generators. Data centers are mission-critical facilities that are designed to operate with high availability and uptime guarantees. As a result, to maintain operation during emergencies such as grid outages, AI data centers require highly reliable backup power sources [10, 45]. Diesel generators are known to emit significant amounts of air pollutants and even hazardous emissions during operation. For example, they emit 200-600 times more NOx than new or controlled existing natural gas-fired power plants for each unit of electricity produced [55]. Nonetheless, there is limited experience with cleaner backup alternatives that can provide comparable reliability in real-world settings, as highlighted by the U.S. Department of Energy in its recent recommendations regarding AI data center infrastructures [6]. Consequently, AI data centers, including those newly built by major technology companies, primarily depend on on-site diesel generators for backup power [6, 10, 45, 56]. For example, in northern Virginia (mostly in Loudoun, Prince William, and Fairfax), the number of permits for data center diesel generators has increased by about 70% since 2023 compared to the total number of permits issued between 2000 and 2022 [56].
@@
While diesel generators need to comply with air quality regulations and typically do not operate over extended periods of time, regular maintenance and testing are essential to ensure their operational reliability. In addition, capacity redundancy is typically followed for diesel generator installations to ensure high availability [58]. Thus, diesel generators represent a major source of on-site air pollutants for data centers and pose a significant health risk to the public [59]. For instance, the total permitted annual emission limits for data centers in northern Virginia are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S. short tons. Assuming that the actual emissions are only 10% of the permitted level, these backup generators could already cause 14,000 asthma symptom cases and 13-19 deaths each year among other health implications, resulting in a total annual public health burden of \$220-300 million throughout the U.S. This includes \$190-260 million in Virginia, West Virginia, Maryland, Pennsylvania, Delaware, New Jersey, New York, and Washington D.C. We show the county-level health cost and the top-10 counties in Figure 1, while deferring the details of calculations to Appendix A.3.
Moreover, due to power grid capacity constraints in many U.S. states, AI data centers are increasingly pressured to vary their loads subject to the grid’s operating conditions, i.e., grid-integrated data centers [60, 6]. This trend may necessitate extended reliance on backup generators, e.g., possibly 15 days per year [6]. Such prolonged usage of diesel generators could substantially elevate AI’s scope-1 air pollution, creating even higher public health costs. Concretely, if the backup generators in northern Virginia emit air pollutants at the maximum permitted level, the total public health cost could reach \$2.2-3.0 billion per year.
What further adds to the public health threat is that many data center generators in a region may operate simultaneously for demand response during grid capacity shortages, potentially resulting in a short-term spike in PM2.5 and NOx emissions that can be particularly harmful [6, 24, 31].
2.2.2 Scope 2
While technology companies have started implementing various initiatives — such as purchasing renewable energy credits and nuclear power from small modular reactors [5, 10, 61] — to lower their (market-based) carbon emissions, the vast majority of U.S. data centers remain directly powered by local power grids with a substantial portion of fossil fuel-based energy sources [10]. Thus, just as AI is accountable for scope-2 carbon emissions, it also contributes to scope-2 air pollution through its electricity usage.

### v2

# The Unpaid Toll: Quantifying and Addressing the Public Health Impact of Data Centers

@@ Abstract
The surging demand for AI has led to a rapid expansion of energy-intensive data centers, impacting the environment through escalating carbon emissions and water consumption. While significant attention has been paid to data centers’ growing environmental footprint, the public health burden, a hidden toll of data centers, has been largely overlooked. Specifically, data centers’ lifecycle, from chip manufacturing to operation, can significantly degrade air quality through emissions of criteria air pollutants such as fine particulate matter, substantially impacting public health. This paper introduces a principled methodology to model lifecycle pollutant emissions for data centers and computing tasks, quantifying the public health impacts. Our findings reveal that training a large AI model comparable to the Llama-3.1 scale can produce air pollutants equivalent to more than 10,000 round trips by car between Los Angeles and New York City. The growing demand for AI is projected to push the total annual public health burden of U.S. data centers up to more than \$20 billion in 2028, rivaling that of on-road emissions of California. Further, the public health costs are more felt in disadvantaged communities, where the per-household health burden could be 200x more than that in less-impacted communities. Finally, we propose a health-informed computing framework that explicitly incorporates public health risk as a key metric for scheduling data center workloads across space and time, which can effectively mitigate adverse health impacts while advancing environmental sustainability. More broadly, we also recommend adopting a standard reporting protocol for the public health impacts of data centers and paying attention to all impacted communities.
@@ 1 Introduction
The rise of artificial intelligence (AI) has numerous potentials to play a transformative role in addressing grand societal challenges, including air quality and public health [1, 2]. For example, by integrating multimodal data from various sources, AI can provide effective tools and actionable insights for pandemic preparedness, disease prevention, healthcare optimization, and air quality management [3, 1]. However, the surging demand for AI — particularly generative AI, as exemplified by the recent popularity of large language models (LLMs) — has driven a rapid increase in computational needs, fueling the unprecedented expansion of energy-intensive data centers. According to the recent Lawrence Berkeley National Lab report [4], AI training and inference are projected to become the dominant workloads and push the U.S. data center electricity consumption to account for 6.7–12.0% of the national total in 2028, up from 4.4% in 2023.
@@
The growing electricity demand of data centers has not only created significant stress on power grid stability [5, 6], but also increasingly impacts the environment through escalating carbon emissions [7, 8] and water consumption [9]. These environmental impacts are driven primarily by the “expansion of AI products and services,” as recently acknowledged by technology companies in their sustainability reports [10]. To mitigate the challenges posed to both power grids and the environment, a range of strategies have been explored, including grid-integrated data centers [6, 11], energy-efficient hardware and software [12, 13, 14], and the adoption of carbon-aware and water-efficient computing practices [15, 9, 16, 17], among others.
@@
The hidden toll of data centers. While the environmental footprint of data centers has garnered attention, the public health burden, a hidden toll of data centers, has been largely overlooked. Across its entire lifecycle — from chip manufacturing to operation — a data center contributes substantially to air quality degradation and public health costs through the emission of various criteria air pollutants. These include fine particulate matter (PM2.5, particles measuring 2.5 micrometers or smaller in diameter that can penetrate deep into lungs and cause serious health effects), sulfur dioxide (SO2), and nitrogen dioxide (NO2). Concretely, the server manufacturing process [18], electricity generation from fossil fuels to power data centers, and the maintenance and usage of diesel backup generators to ensure continuous data center operation all produce significant amounts of criteria air pollutants. Moreover, the distinct spatial-temporal heterogeneities of emission sources suggest that focusing solely on reducing data centers’ carbon footprints may not minimize its emissions of criteria air pollutants or the resulting public health impacts (Section 6.2).
@@
Exposure to criteria air pollutants is directly and causally linked to various adverse health outcomes, [^2] While we focus on public health, we note that the impacts of criteria air pollutants extend beyond humans and include harms to environmentally sensitive areas, such as some national parks and wilderness areas which, classified as “Class 1 areas” under the Clean Air Act, require special air protection [19]. including premature mortality, lung cancer, asthma, heart attacks, cardiovascular diseases, and even cognitive decline, especially for the elderly and vulnerable individuals with pre-existing conditions [20, 21, 22]. Moreover, even short-term (hours to days) PM2.5 exposure is harmful and deadly, accounting for approximately 1 million premature deaths per year from 2000 to 2019 and representing 2% of total global deaths [23].
Criteria air pollutants are not confined to the immediate vicinity of their emission sources; they can travel hundreds of miles through a dispersion process (i.e., cross-state air pollution) [24, 25], impacting public health across vast regions. Further, PM2.5 is considered “non-threshold,” i.e., there is no absolutely safe exposure level [26]. Thus, compliance with the national/regional air quality standards does not necessarily ensure the air is healthy.
@@
Importantly, along with transportation and industrial activities, electricity generation is a key contributor to ambient air pollution with substantial public health impacts [28, 30, 31]. For example, a recent study [32] shows that, between 1999 and 2020, a total of 460,000 excess deaths were attributed to PM2.5 generated by coal-fired power plants alone in the U.S. As highlighted by the U.S. EPA [30], despite years of progress, power plants “remain a leading source of air, water, and land pollution that affects communities nationwide.” In Europe, the public health cost of air pollution from power plants is valued at approximately 1% of the gross domestic product (GDP), according to the European Environment Agency’s study in 2024 [33].
The public health outcomes of data centers due to their emission of criteria air pollutants lead to various losses, such as hospitalizations, medication usage, emergency room visits, school loss days, and lost workdays. Nonetheless, despite recent policy efforts [34, 35], the tangible and growing public health impacts of data centers have remained under the radar, almost entirely omitted from today’s risk assessments and sustainability reports [36, 37, 10].
@@
Quantifying and addressing the public health impacts of data centers. In this paper, we uncover and quantify the hidden public health impacts of data centers. We introduce a principled methodology to model the emission of criteria air pollutants associated with a computing task and data center across three distinct scopes: emissions from the maintenance and operation of backup generators (Scope 1), emissions from fossil fuel combustion for electricity generation (Scope 2), and emissions resulting from the manufacturing of server hardware (Scope 3). Then, we analyze the dispersion of criteria air pollutants and the resulting public health impacts.
@@
As the U.S. hosts nearly half of the world’s data centers [38] and the EPA data excludes other regions [39], our empirical study focuses on the 48 contiguous U.S. states plus Washington D.C. [^3] If located in countries with higher population densities, more pollutant-intensive electricity mixes, or less stringent air quality standards, the same data centers would likely lead to more premature deaths and other adverse health impacts than in the U.S. We recommend further research on the public health impact of non-U.S. data centers. Our main results (Section 5) focus on the scope-1 and scope-2 health impacts of U.S. data centers and, specifically, LLM training. Using the reduced-complexity modeling tool COBRA (CO-Benefits Risk Assessment) provided by the EPA [39], our analysis demonstrates that driven by the growing demand for AI, the U.S. data centers could contribute to, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths in 2028, exceeding 1/3 of asthma deaths in the U.S. each year [40]. The overall public health costs could reach more than \$20 billion, rival or even top those of on-road emissions of the largest U.S. states such as California with $\sim$35 million registered vehicles [41]. Moreover, depending on the location, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a car for more than 10,000 round trips between Los Angeles and New York City (LA-NYC), resulting in a health cost that even exceeds 120% of the training electricity cost.
Importantly, although the public health impact of data centers may be modest at the national level, it is geographically concentrated, with certain regions and communities bearing a disproportionate share. In particular, some low-income counties experience significantly greater health costs, with per-household burdens exceeding those in other counties by more than 200-fold.
@@
Furthermore, to highlight scope-1 health impacts, we examine data center backup generators in Virginia, which hosts one of the largest concentrations of data centers in the world [42]. Our analysis shows that, assuming the actual emissions are only 10% of the permitted level based on the historical reports and future projections [42, 43, 44], the data center backup generators registered in Virginia (mostly in Loudoun, Prince William, and Fairfax) could already cause 14,000 asthma symptom cases among other health outcomes and a total public health burden of \$220-300 million per year, impacting residents in multiple surrounding states and as far as Florida (Section 3.1). If these data centers emit air pollutants at the maximum permitted level, the total public health cost will become 10-fold and reach \$2.2-3.0 billion per year.
To address the growing and uneven distribution of health burdens, we propose health-informed computing that explicitly incorporates the public health risk as a key metric when scheduling data center workloads. Specifically, by exploiting the spatial-temporal variations of public health impacts and using spatial load shifting as a case study, we demonstrate that the health-informed approach can significantly reduce the health cost compared to the baseline, while continuing to offer meaningful electricity cost savings and reductions in carbon emissions.
@@
Finally, we provide broader recommendations to address the increasing public health impact of data centers (Section 7). We recommend technology companies adopt a standard reporting protocol for criteria air pollutants and public health impacts in their AI model cards and sustainability reports and pay attention to all impacted communities.
@@
To summarize, wile AI and data centers offer many societal benefits, our study sheds light on, quantifies, and addresses the often overlooked negative externalities of their resource demand, particularly the public health impact. We also urge further research to comprehensively address the public health implications when developing data centers in the future, ensuring that the growth of data centers does not exacerbate the health burden.
Disclaimer. The results presented in this paper are not intended to encourage or discourage the construction of data centers, nor should they be used to support or oppose any specific project, which requires more detailed and context-specific evaluation. We do not take a position on decisions related to any specific data centers or the use of AI, but instead provide a quantitative assessment of the potential public health impacts of the data center industry.
2 Background on Air Pollutants
This section provides background on criteria air pollutants and U.S. air quality policies. Other countries have similar policies in place to safeguard public health, although their levels of enforcement strictness often differ [33].
@@ 2 Background on Air Pollutants
While the U.S. has generally better air quality than many other countries, 4 in 10 people in the U.S. still live with unhealthy levels of air pollution, according to the “State of the Air 2024” report published by the American Lung Association [50]. In 2019 (the latest year of data provided by the World Health Organization, or WHO, as of November 2024), an estimate of 93,886 deaths in the U.S. were attributed to ambient air pollution [51]. In fact, the EPA’s recently tightened standard for PM2.5 sets an annual average limit of 9 $\mu g/m^{3}$, considerably higher than the WHO’s recommended level of 5 $\mu g/m^{3}$ [52, 48]. Moreover, the EPA projects that 53 U.S. counties, including 23 in the already most populous state of California, would fail to meet the revised national annual PM2.5 standard in 2032 [53].
@@
Although CO2 is broadly classified by the EPA as an air pollutant following the U.S. Supreme Court ruling in 2007 [54] and contributes to long-term climate change, it often does not cause the same immediate health impacts as criteria pollutants [55]. In the U.S., CO2 and other greenhouse gases are subject to different EPA regulations from those for criteria air pollutants. Thus, for the sake of presentation in this paper, we use “air pollutants” to solely refer to criteria air pollutants wherever applicable.
3 Data Centers’ Contribution to Air Pollutants
Figure 1: The overview of data centers’ contribution to air pollutants and public health impacts. Scope-1 and scope-2 impacts occur during the operation of data centers (“operational”), whereas scope-3 impacts arise from activities across the supply chain (“embodied”).
This section presents an overview of data centers’ impact on air quality and contribution to criteria air pollutants throughout its lifecycle across three scopes (Fig. 1). The scoping definition in this paper parallels the well-established greenhouse gas protocol [56]. Specifically, scope-1 and scope-2 air pollutants primarily originate from onsite generators and power plants, collectively referred to as operational emissions, while scope-3 pollutants arise from the supply chain and are referred to as embodied emissions.
3.1 Scope 1: Onsite Generator
While the construction phase of a data center directly increases air pollutant emissions, its amortized health impact over a typical 15–20 year lifespan is negligible. Therefore, we focus on onsite backup generators as the primary source of scope-1 direct air pollutants.
@@ 3.1 Scope 1: Onsite Generator
Data centers are mission-critical facilities that are designed to operate with high availability and uptime guarantees. As a result, to maintain operation during emergencies such as grid outages, data centers require highly reliable backup power sources [10, 37]. Diesel generators are known to emit significant amounts of air pollutants and even hazardous emissions during operation [57]. For example, they emit 200-600 times more NOx than new or controlled existing natural gas-fired power plants for each unit of electricity produced [58]. Moreover, capacity redundancy is typically followed for diesel generator installations to ensure high availability [59]. Nonetheless, there is limited practical experience with cleaner backup alternatives that can provide comparable reliability at scale in real-world settings in the near term, as highlighted by the U.S. Department of Energy in its recent recommendations regarding AI data center infrastructures [6].
Consequently, the vast majority of data centers, even including those newly built by major technology companies, depend on onsite diesel generators for backup power [6, 60]. For example, in Northern Virginia (mostly in Loudoun, Prince William, and Fairfax), the number of permits for data center diesel generators has increased by more than 70% since 2023 compared to the total number of permits issued between 2000 and 2022 [60]. Importantly, nearly all the diesel generators are Tier 2, which have significantly higher emission rates than Tier 4 units [61, 60]. The total permitted annual emission limits for these diesel generators are approximately 13,000 tons of NOx, 1,400 tons of VOCs, 50 tons of SO2, and 600 tons of PM2.5, all in U.S. short tons.
@@
While diesel generators need to comply with air quality regulations and typically do not operate over extended periods of time, regular maintenance and testing are essential to ensure their operational reliability. A recent report by the state of Virginia [42] found that the actual air pollutant emissions from backup generators at Virginia’s data centers reached approximately 7% of the total permitted amounts in 2023, primarily for maintenance. Likewise, the actual emissions took up 3% to 12% of the permitted levels for some data centers in Quincy, Washington [43].
Moreover, the U.S. EPA recently issued a clarification that would allow data centers to run backup generators for up to 50 hours a year (or roughly 10% of the permits that typically allow 500 hours per year) to participate in demand response—a program designed to reduce grid demand during peak hours, which is increasingly activated as surging data center demand strains grid capacity [44]. This trend may necessitate extended reliance on backup generators [6]. What further adds to the public health impact is that many data center generators in a region may operate simultaneously for demand response during grid capacity shortages, potentially resulting in a short-term spike in PM2.5 and NOx emissions that can be particularly harmful [6, 23, 48]. For example, from June 23 to 25, 2025, some data centers in Loudoun County, Virginia, were instructed to run their on-site diesel generators for demand response, releasing large amounts of air pollutants and “black smoke” [61].
@@
The high emission rate from onsite generators, combined with extended operation for maintenance and demand response beyond grid outages, could pose serious health risks, especially in regions with a concentration of large data centers. To illustrate this point, we consider the data centers’ onsite generators in Virginia. Assuming that the actual emissions are 10% of the permitted level as a reference case that reflects both the historical reports and future demand response projections [42, 43, 44],  [^4] If the actual percentage is xx%, our value will be approximately scaled by $\frac{x}{10}$ [22]. the backup generators could already cause 14,000 asthma symptom cases and 13-19 deaths each year among other health implications, resulting in a total annual public health burden of \$220-300 million throughout the U.S. This includes \$190-260 million in Virginia, West Virginia, Maryland, Pennsylvania, Delaware, New Jersey, New York, and Washington D.C. We show the county-level health cost and the top-10 counties in Figure 2, while deferring the details of calculations to Appendix A.1. If the diesel generators in Northern Virginia emit air pollutants at the maximum permitted level, the emission of NOx could even exceed half of the annual total emissions by all sources in the region [42], resulting in a total public health cost of \$2.2-3.0 billion per year.
3.2 Scope 2: Power Plants
Just as data centers are accountable for scope-2 carbon emissions, they also contribute to scope-2 air pollution through their electricity usage.